simplify 1/2 of the square root 28 and the squareroot of 4000x squared y
Answer by stanbon(75887)

simplify 1/2 of the square root 28
= (1/2)sqrt(28)
= (1/2)*2Sqrt(7)
= sqrt(7)
--------------
simplify the squareroot of 4000x squared y
= sqrt(4000x^2y)
= sqrt(400x^2*10y)
= 20x*sqrt(10y)
===================
Cheers,
Stan H.